OpenStudy (mrchair):

Solve for m. -3 + m 9 = 10 -30 63 87 93

OpenStudy (mrchair):

-3+m/9=10

OpenStudy (mrchair):

@ARMYRANGERS

OpenStudy (armyrangers):

What do you think it is

OpenStudy (mrchair):

-30

OpenStudy (mrchair):

m=-30

OpenStudy (armyrangers):

no you have to do 10*9 and that gives you what

OpenStudy (mrchair):

90

OpenStudy (mrchair):

so that gives me 87?

OpenStudy (armyrangers):

No it is negative 3 so you have to add 3 to make it 0

sammixboo (sammixboo):

\[(-3)+m/9 = 10\] \[(-3)+m/9*9= 10*9\] \[(-3)+m= 90\] \[(-3) - (-3)+m= 90 - (-3)\] \(m= 90 - (-3)\) What would your final answer would be?

OpenStudy (armyrangers):

Why did you take my answer I WAs getting there I was trying to help him learn it

sammixboo (sammixboo):

Remember when subtracting integers that have the opposite sign, you change the SUBTRACTION sign into an ADDITION sign and change the sign of the integer to the right, which is\(\ -3\) , to the opposite sign, so now \(-3\) is \(+3\). Now the equation looks like this: \(90 + 3\) Add those up and there is your answer.

OpenStudy (mrchair):

thank you guys